(a) If a 500-mL glass beaker is filled to the brim with ethyl alcohol at a temperature of 5.00ºC, how much will overflow when its temperature reaches 22.0ºC? The coefficient of volume expansion for glass is 27x10-6 and for ethyl alcohol is 1100x10-6. (b) How much water would overflow under the same conditions? The coefficient of volume expansion for water is 210x10-6.
Answer #1

AV (a) AV= (Falcohol - Bg )x VX AT = (1100 - 27)x10-6x soo x (22-5) = 9.12 m2 Answer (b) AV= (210 - 27)x105°x soo x (22-5) =